a) How many arrangements are there of the word “GRANDMASTER”?

b) Suppose a lottery consists of picking 6 different numbers from 1 to 52. (The order you pick them doesn’t matter). How many possible lottery tickets are there?
Answer #1

the word N = IL (a) Mumba GRAND MASTER of corrangements of Total number letters in the given coord is Repetition letter A mi)cond iet is geven 03 , T- N 52. 2C6 cd T= 521 in NI t 6! (52-6)! کیل Y!(N-Y)! T= 52! 6 !X 46! T= 52X SIX SOX 494 48 847 68 5X